# Empowering Biomedical Discovery with AI Agents

### v1

@@ 5.4 Reasoning modules
Reasoning with feedback. Experimental and human feedback can help AI agents to improve reasoning and planning processes [12, 69, 161]. This feedback may include agent-human interaction (Section 5.2) and responses from agents, which can be complementary biological assays quantifying downstream effects of target molecules [182]. In each reasoning cycle, React [12] incorporates insights from previous actions to refine its thought process and inform future actions. LLM-Planner [183] dynamically adjusts plans based on new observations in an embodied environment. Inner Monologue [138] uses both passive and active scene descriptions and feedback from recent actions to guide future actions. Voyager [69] improves planning for subsequent steps by considering environment feedback, execution errors, and self-verification.
Beyond external feedback, an agent’s feedback mechanism enables self-assessing the initial plan [182, 184]. Techniques like self-refine [182] revise action outputs based on the LLM evaluation, the self-check [182] mechanism allows the agent to review and adjust its reasoning, and reflection [13] mechanisms use prompt agents to update their decision-making. These techniques incorporate feedback from biologists, such as exploring experimental methods and environmental constraints like lab inventory (Figure 5d).
6 Challenges
The perspective outlines key steps to implement AI agents in biomedical research and highlights areas that can benefit from agentic AI. Challenges remain and may, in some cases, be amplified when multi-agent systems become available. Below are technical obstacles and challenges to overcome (Figure 6).
6.1 Robustness and reliability
A barrier facing the deployment of agent systems – specifically those categorized within Levels 2 and 3 as discussed in Section 2 – is their propensity for generating unreliable predictions, including the hallucination of non-factual information, reasoning errors, systematic biases, and failures in planning when connected with tools and experimental platforms. These issues can be exacerbated by overconfidence in such flawed predictions (agents are unaware of their knowledge gaps) and high sensitivity to the precise formulation of queries, particularly in the context of LLM-based agents.
Such behavior has been traced to the pretraining phase of these models. By contrasting the predicted word sequence and the sequence present in the training data, the autoregressive loss function influences the subsequent model performance with the probability distribution of its inputs, the sequence of generated outputs, and the frequency of tasks it encounters during training [185]. As a result, model performance degrades on task variants that deviate from the assumptions made during training [186].
@@ 6.1 Robustness and reliability
An overarching concern is that advanced capabilities come at the cost of compromised transparency and the risk of misalignment. For instance, integrating human feedback can promote desirable agent behavior, but it can also exacerbate persuasive abilities, echoing false beliefs [193]. Furthermore, fine-tuning existing models with new data can compromise their original alignment, challenging the integrity of the AI agent’s intended purpose [194]. Jailbreak attacks can similarly affect post-deployment, highlighting the need for rigorous evaluation [195].
6.2 Evaluation protocols
@@
With more AI agents being developed, frameworks for biologists and lay user evaluations need to assess axes of agent performance beyond accuracy. Evaluating AI agents requires an analysis of their theoretical capabilities and an assessment of practical implications, including ethical considerations, regulatory compliance, and the ability to integrate into discovery workflows. The challenge lies in developing evaluations that consider these diverse factors. Agents that integrate ML tools, particularly those developed by corporations, may undergo updates without prior notice to users. This poses challenges for reproducibility, as updates may alter the model’s behavior or performance without researchers being aware. The scientific community requires transparent change logs and version control for agents, akin to practice in software development.
@@
Evaluation frameworks consider either a holistic evaluation of LLMs [196, 197] or weak spots such as task framing [198, 199], long temporal dependencies, invalid formatting or refusal to follow instructions [200]. A caveat of such methods is the risk of evaluating how well the agents have learned to use specific APIs versus general results grounded in real-world interaction. Another challenge in evaluating agents is that biological systems are inherently dynamic, characterized by non-stationary distributions that evolve due to genetic mutations, environmental changes, and evolutionary pressures. Agents trained on static datasets may struggle to accurately model or predict outcomes in these changing systems. The challenge lies in developing agents capable of adapting to or continuously learning from new data, ensuring their predictions remain accurate as the underlying biological systems change. Techniques such as online learning, transfer learning, and reinforcement learning can be used to address this issue, but they come with their own set of challenges related to data availability and model complexity.
6.3 Dataset generation
@@
As laid out, the vision of advanced agents in biology requires the capability of seeking, aggregating, perceiving, and reasoning over data from various modalities, created using differing specifications and with inherent variation in quality and volume. To support this vision, there is a critical need for large, open datasets that are both comprehensive and accessible, enabling the development of models across biological applications. Much human effort in building complex systems for biology is dedicated to gathering and preparing such data for use in ML models (e.g., specific to a particular modality, such as graphs, time series, or discrete sequences [201]). This requires vetting processes and clear criteria for assessing the reliability and applicability of datasets.
@@ 6.4 Governance of AI agents
Safe adoption of AI agents requires addressing concerns of safe deployment. Aligning ML tools, such as LLMs, with ethical standards remains an open challenge, and ensuring the alignment of the agent as a digital entity raises complexity. Guidelines concerning human-agent interactions are underdeveloped despite the potential for both unintended harmful consequences and malicious intent. Tang et al. [85] describe a safeguarding framework that includes training, licensing, and mandatory safety and ethical compliance checks for agents.
As AI agents become more integral to workflows in biological domains, monitoring their behavior grows increasingly complex. Currently, verifying the accuracy and trustworthiness of agent outputs is not straightforward, with only a limited number of systems capable of linking generated content to relevant references. Assessing the synthesized knowledge may be impractical and unattainable as agents evolve further. When agents’ capabilities become comparable to those of human experts, the risk of becoming overly reliant on AI increases, which could lead to a decrease in human expertise. In the worst-case scenario, such reliance could introduce a broad spectrum of safety hazards due to inadequate oversight.
6.5 Risks and safeguards
Developing autonomous experiments and deployment that do not include careful planning, broad consultation, competent execution, and ongoing adaptation might create long-term harms that outweigh the benefits. Although anticipating all potential complications is impossible, exploring possible problems early and frequently could reduce the expected cost of such issues. The space of ethical considerations relevant to AI agents is too broad to canvass comprehensively here, but this section highlights a few key categories.
@@ 6.5 Risks and safeguards
LLM agents leverage the world knowledge and general reasoning abilities of LLMs obtained during pretraining for solving robotics and planning tasks. However, while considerable effort has been made to teach the robots the “dos,” the “don’ts” received less attention. In agent systems, teaching the robots the “don’ts” is crucial: conveying instructions about prohibited actions, assessing the robot’s comprehension of these restrictions, and ensuring compliance [204]. For LLM agents, Yang et al. [204] developed a plug-in safety chip, a queryable safety constraint module that translates natural language constraints into formal safety constraints that the robot adheres to. Experiments with robots highlight the potential for integrating formal methods with LLMs for robotic control.
@@
LLMs trained in code completion can write Python programs from docstrings [205]. Given natural language commands, these code-writing LLMs can be re-purposed to write robot policy code. However, if the translation inaccurately reflects the intended safety constraints, it could lead to either overly restrictive behavior, preventing the robot from performing its tasks effectively, or insufficiently stringent constraints, leading to safety violations. Writing robot policy code entails using LLMs trained on code completion to write the code from natural language commands [206]. However, the robot policy code is less reliable for enforcing safety constraints than verifiable safe operations that satisfy standards such as ISO 61508. The approach assumes that all given instructions are feasible and lacks a mechanism to predict the correctness of a response before execution. However, due to their reliance on patterns in the training data, LLMs might generate syntactically correct but semantically inappropriate code. They also struggle with understanding physical constraints of the environment in which robots operate. Additionally, generalizing plans across robotic embodiments is brittle with current LLMs. This implies a limitation in what can be generalized, particularly without extensive data collection and adaptability of the generated policies.
7 Outlook
@@
Biomedical research is undergoing a transformative era with advances in computational intelligence. Presently, AI’s role is constrained to assistive tools in low-stake and narrow tasks where scientists can review the results. We outline agent-based AI to pave the way for systems capable of skeptical learning and reasoning that consist of LLM-based systems and other ML tools, experimental platforms, humans, or even combinations of them. The continual nature of human-AI interaction creates a path to achieve this vision once focused on preventing and learning from mistakes. Building trustworthy sandboxes [207], where AI agents can fail and learn from their mistakes, is one way to achieve this. This involves developing AI agents that perform tasks and consider the boundary of their generalization ability, fostering natural and artificial intelligence.
@@ Acknowledgments
We gratefully acknowledge the support of NIH R01-HD108794, NSF CAREER 2339524, US DoD FA8702-15-D-0001, awards from Harvard Data Science Initiative, Amazon Faculty Research, Google Research Scholar Program, AstraZeneca Research, Roche Alliance with Distinguished Scientists, Sanofi iDEA-iTECH Award, Pfizer Research, Chan Zuckerberg Initiative, John and Virginia Kaneb Fellowship award at Harvard Medical School, Aligning Science Across Parkinson’s (ASAP) Initiative, Biswas Computational Biology Initiative in partnership with the Milken Institute, and Kempner Institute for the Study of Natural and Artificial Intelligence at Harvard University. A.F. is supported by the Kempner Institute Graduate Fellowship. V.G. is supported by the Medical Research Council, MR/W00710X/1. Y.E. is supported by grant T32 HG002295 from the National Human Genome Research Institute and the NSDEG fellowship. The authors would like to thank Owen Queen, Alejandro Velez-Arce, and Ruth Johnson for their constructive comments on the draft manuscript. Any opinions, findings, conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the funders.
@@
Figure 1: Empowering biomedical research with AI agents. AI agents pave the way for "AI scientists" capable of skeptical learning and reasoning. These multi-agent systems consist of agents based on conversable large language models (LLMs) and can coordinate machine learning (ML) tools, experimental platforms, humans, or even combinations of them. Robotic agent, AI agent that operates robotic hardware for physical experiments; Database agent, AI agent that can information in databases via ‘function calling’ and APIs; Reasoning agent, AI agent capable of direct reasoning and reasoning with feedback; Hypothesis agent, AI agent that is creative and skeptical when developing hypotheses, capable of characterizing its own uncertainty and using that as a driver to refine its scientific knowledge bases; Brainstorming agent, AI agent that generates a broad spectrum of research ideas; Search engine agent, AI agent that uses search engines as tools to rapidly gather information; Analysis agent, AI agent capable of analyzing experimental results to summarize findings and synthesize concepts; Experimental planning agent, AI agent that optimizes an experimental protocol for execution.
@@
Figure 2: Evolving use of data-driven models in research. Data-driven approaches, from databases and search engines, machine learning, and interactive learning models to advanced agent systems (Section 2), have reshaped biomedical research throughout the last several decades.
@@
Figure 3: Diverse configurations of AI agents in biology – from an LLM-based AI agent to a multi-agent system with AI models, tools, and integrated physical devices. a. By programming an LLM with the role, one LLM-based agent, equipped with memory and reasoning abilities, performs multi-modal perception and utilizes a range of tools, e.g., web lab tools, to accomplish specified tasks. b-e. Leveraging AI agents equipped with diverse roles, perception modules, tools, and domain knowledge enables collaboration between agents and scientists. This collaboration can adopt various schemes, such as expert consultation, debate, brainstorming, and roundtable discussions. f. Multi-agent systems can establish a self-driving laboratory wherein numerous agents collaborate on multiple iterations of biological research assisted by humans. Each cycle of research encompasses the generation of hypotheses, the design of experiments, the execution of experiments both in silico and in vitro, and the analysis of results. Computing agent, AI agent that utilizes computational models as tools; Decision agent, AI agent that makes decisions in response to given conditions; Database agent, AI agent that retrieves relevant information from databases; Reasoning agent, AI agent capable of direct reasoning and reasoning with feedback; Expert agent, AI agent that provides professional consultation based on reliable sources, such as domain expertise, feedback from human experts, and the results of specific tools. Hypothesis agent, AI agent capable of skeptical learning and reasoning to generate hypotheses; Planner agent, AI agent that devises plans for future actions; In silico/vitro agent, AI agent that uses tools in silico or in vitro environment.
@@
Figure 4: AI agents consist of four key modules: perception, interaction, reasoning, and memory modules. Perception interprets multi-modal environmental data. Interaction facilitates engagement with the environment, encompassing human-agent interactions, multi-agent interactions, and tool use. Memory is responsible for the storage and retrieval of knowledge, while Learning focuses on the acquisition and updating of knowledge. Reasoning, with or without environmental feedback, plays a crucial role in planning and decision-making processes.
@@
Figure 5: Components of biomedical AI agents. a. Use of a short-term memory module to recall previous relevant experiments for small molecule inhibitor design. b. Use of a long-term memory module to retrieve relevant information for target selection for a disease. c. Use of reasoning without scientist feedback in gene prioritization for phenotype analysis. d. Use of reasoning with feedback from scientists to select an alternative experimental approach.
@@
Figure 6: Challenges for AI agents in scientific discovery. a. Shown are critical challenges–including robustness and reliability, evaluation protocols, dataset generation, governance, and risks—alongside b. strategic approaches to address them.
Autonomy levels Scientific discovery Scientist-AI agent roles
Hypothesis Experiment Reasoning
@@
In-context learning Ability of models, such as LLMs, to achieve new tasks based on the context provided within contextual prompts, without requiring explicit model training

### v2

@@ Introduction
Reasoning with feedback. Experimental and human feedback can help AI agents to improve reasoning and planning processes [11, 68, 149]. This feedback may include agent-human interaction and responses from agents, which can be complementary biological assays quantifying downstream effects of target molecules [170]. In each reasoning cycle, React [11] incorporates insights from previous actions to refine its thought process and inform future actions. LLM-Planner [171] dynamically adjusts plans based on new observations in an embodied environment. Inner Monologue [126] uses both passive and active scene descriptions and feedback from recent actions to guide future actions. Voyager [68] improves planning for subsequent steps by considering environment feedback, execution errors, and self-verification.
Beyond external feedback, an agent’s feedback mechanism enables self-assessing the initial plan [170, 172]. Techniques like self-refine [170] revise action outputs based on the LLM evaluation, the self-check [170] mechanism allows the agent to review and adjust its reasoning, and reflection [12] mechanisms use prompt agents to update their decision-making. These techniques incorporate feedback from biologists, such as exploring experimental methods and environmental constraints like lab inventory (Figure 5d). Reasoning capabilities are necessary for generating hypotheses and conducting experiments. Generating novel hypotheses requires modeling general biomedical knowledge, the specific information on the current state of a biological system, and consideration of potential next steps. LLM-based agents can generate hypotheses through in-context reasoning, but careful selection is necessary to ensure high-quality hypotheses [173].
Challenges
The perspective outlines key steps to implement AI agents in biomedical research and highlights areas that can benefit from agentic AI. Challenges remain and may, in some cases, be amplified when multi-agent systems become available (Figure 6).
Robustness and reliability
A barrier facing the deployment of agent systems – specifically those categorized within Levels 2 and 3 as discussed inTable 1 – is their propensity for generating unreliable predictions, including the hallucination of non-factual information, reasoning errors, systematic biases, and failures in planning when connected with tools and experimental platforms. These issues can be exacerbated by overconfidence in such flawed predictions (agents lack awareness of their knowledge gaps) and high sensitivity to the precise formulation of queries, particularly in the context of LLM-based agents. This behavior has been traced to how these models are trained. In particular, autoregressive loss compares the predicted word sequence with the actual sequence in the training data. The performance of a model trained with this loss is determined by three factors: the probability distribution of the inputs, the sequence of generated outputs, and the frequency of different tasks encountered during training [174]. As a result, model performance degrades on task variants that deviate from the assumptions made during training [175].
@@
One concern is that advanced capabilities come at the cost of compromised transparency and the risk of misalignment. For instance, integrating human feedback can promote desirable agent behavior, but it can also exacerbate persuasive abilities, echoing false beliefs [182]. Fine-tuning existing models with new data can compromise their original alignment, challenging the integrity of the AI agent’s intended purpose [183]. Jailbreak attacks can similarly affect post-deployment, highlighting the need for rigorous evaluation [184].
Errors are inevitable in complex multi-agent systems, making their management crucial to maintaining system robustness and reliability. Due to their interactive nature, these systems are sensitive to compounding errors, where small issues can escalate into significant problems if not addressed promptly. Effective error management strategies are essential for diagnosing, localizing, and mitigating such errors.
Evaluation protocols
@@
With more AI agents being developed, frameworks for biologists and lay user evaluations need to assess axes of agent performance beyond accuracy. Evaluating AI agents requires an analysis of their theoretical capabilities and an assessment of practical implications, including ethical considerations, regulatory compliance, and the ability to integrate into discovery workflows. The challenge lies in developing evaluations that consider these diverse factors. Agents that integrate ML tools, particularly those developed by corporations, may undergo updates without prior notice to users. This poses challenges for reproducibility, as updates may alter the model’s behavior or performance without researchers being aware. The scientific community needs transparent change logs and version control for agents, akin to practice in software development.
@@
Existing evaluation frameworks consider either holistic evaluations [185, 186] or benchmark the models for weak spots such as task framing [187, 188], long temporal dependencies, invalid formatting or refusal to follow instructions [189]. A caveat of such frameworks is the risk of evaluating how well the agents have learned to use specific APIs versus general results grounded in real-world interaction. Another challenge in evaluating agents is that biological systems are inherently dynamic, characterized by non-stationary distributions that evolve due to genetic mutations, environmental changes, and evolutionary pressures. Agents trained on static datasets may struggle to accurately model or predict outcomes in these changing systems. The challenge lies in developing agents capable of adapting to or continuously learning from new data, ensuring their predictions remain accurate as the underlying biological systems change. Techniques such as online learning, transfer learning, and reinforcement learning can be used to address this issue, but they come with their own set of challenges related to data availability and model complexity. Another challenge is the lack of standardization in biomedical discovery workflows, including data generation protocols that vary based on factors like disease cell lines, dosage levels, and time points [190]. This variability complicates the evaluation of agents for experimental planning. Evaluation of agents that use computational tools and databases will benefit from the increasing availability of standardized and application APIs [191, 192].
Dataset generation
@@
As laid out, the vision for biomedical AI agents requires the capability of seeking, aggregating, perceiving, and reasoning over data from various modalities, created using differing specifications and with inherent variation in quality and volume. To support this vision, there is a critical need for large, open datasets that are both comprehensive and accessible, enabling the development of models across biological applications. Much human effort in building systems for biomedical research is dedicated to gathering and preparing such data for use in ML models (e.g., specific to a particular modality, such as graphs, time series, or discrete sequences [193]). This requires vetting processes and clear criteria for assessing the reliability and applicability of datasets.
@@
Safe adoption of AI agents requires addressing concerns of safe deployment. Aligning ML tools, such as LLMs, with ethical standards remains an open challenge, and ensuring the alignment of the agent as a digital entity raises complexity. Guidelines concerning human-agent interactions are underdeveloped despite the potential for unintended harmful consequences and malicious intent. Safeguarding frameworks are developed that include training, licensing, and mandatory safety and ethical compliance checks for agents [86].
As AI agents become more integral to workflows in biological domains, monitoring their behavior grows increasingly complex. Currently, verifying the accuracy and trustworthiness of agent outputs is not straightforward, with only a limited number of systems capable of linking generated content to relevant references. It is essential to develop robust verification systems that can provide traceable references for generated content. Assessing the synthesized knowledge may be impractical and unattainable as agents evolve further. When agents’ capabilities become comparable to those of human experts, the risk of becoming overly reliant on AI increases, which could lead to a decrease in human expertise. In the worst-case scenario, such reliance could introduce a broad spectrum of safety hazards due to inadequate oversight. To address these challenges, human-in-the-loop approaches can help maintain accountability. Continuous training and development of human expertise alongside AI can mitigate the risks of over-reliance on AI.
Risks and safeguards
Autonomous experiments that do not include careful planning, broad consultation, competent execution, and ongoing adaptation might create long-term harms that outweigh the benefits. Although anticipating all potential complications is impossible, exploring possible problems early and frequently could reduce the expected cost of such issues. The ethical and technical considerations relevant to AI agents are vast and deeply interconnected, particularly in biomedicine. This section will highlight some key categories.
@@
Agents leverage LLMs’ world knowledge and general reasoning abilities obtained during pretraining for robotics and planning. However, while efforts have been made to teach the robots the “dos,” the “don’ts” received less attention. Teaching robot agents the “don’ts” is crucial to convey instructions about prohibited actions, assessing the agent’s understanding of these restrictions, and ensuring compliance [196]. For LLM agents, plug-in safety chips [196] feature safety constraint modules that translate natural language constraints into formal safety constraints for the robot to adhere to. Experiments with robots highlight the potential for integrating formal methods with LLMs for robotic control.
@@
LLMs trained in code completion can write Python programs from docstrings [197] by training the model on the code completion task to write the code based on natural language commands [198]. Given natural language commands, these code-writing LLMs can be re-purposed to write robot policy code. However, if the translation inaccurately reflects the intended safety constraints, it could lead to either overly restrictive behavior, preventing the robot from performing its tasks effectively, or insufficiently stringent constraints, leading to safety violations. However, the robot policy code is less reliable for enforcing safety constraints than verifiable safe operations that satisfy standards such as ISO 61508. The approach assumes that all given instructions are feasible and lacks a mechanism to predict the correctness of a response before execution. However, due to their reliance on patterns in the training data, LLMs might generate syntactically correct but semantically inappropriate code. Additionally, generalizing plans across robotic embodiments is brittle with current LLMs.
Addressing the ethical implications of AI agents is paramount, given the direct impact on human and animal health and life. The handling of sensitive biological and medical data necessitates robust technological and regulatory measures to ensure security and confidentiality. One promising approach involves using privacy-preserving computation to train agents to protect the privacy of highly sensitive medical data. Homomorphic encryption can secure sensitive data by allowing computations on encrypted data and federated learning techniques allow training agents in a distributed manner without the need to centralize from across sites into a single data repository.
Algorithmic fairness is equally crucial, as biased AI agents can exacerbate health disparities across patients and increase inequalities in the volume of generated datasets and quality of biomedical knowledge, especially for diseases in long-tailed distributions in biological systems. The development of techniques such as adversarial debiasing and fair representation learning offers promising avenues to mitigate these risks. In addition, the black-box nature of these compound AI systems poses another challenge, particularly in healthcare, where interpretability is vital for clinical adoption and patient trust. To provide clearer rationales for the agents’ decisions and make them more acceptable to users, it will become crucial to incorporate interactive dialogue systems that explain agentic outputs through natural language conversations. Ethical considerations surrounding biosafety emerge as AI agents advance toward Level 3 agents. These issues intersect with ongoing debates in bioethics regarding synthetic biology, artificial organisms, and AI-driven life forms, requiring regulatory guidance and engagement from bioethicists and safety experts to ensure alignment with societal values and safety standards.
Challenges uniquely relevant for biomedical AI agents
Biomedical AI agents face several unique challenges that distinguish them from other applications of AI. While strong AI agents have the potential to mitigate some of these challenges, their implementation in biomedical research requires careful consideration. One of the primary challenges is the need for robust and reliable systems capable of reasoning, planning, and executing actions in both virtual and hybrid virtual-physical environments. For instance, natural language reasoning chains can enhance the interpretability of an agent’s actions and contextual outcomes, aiding researchers in understanding AI-generated insights. However, certain challenges persist that can delay the reliable implementation of AI agents or even cause harm if these systems are deployed prematurely. A critical issue is the difficulty in distinguishing between correlation and causality. Current AI agents often struggle with generating strong hypotheses, reasoning, and conducting experimental validations, tasks that typically require advanced AI systems (Level 3 agents) or human intervention. Moreover, AI agents need improved interfaces to interact safely and effectively with high-throughput experimental platforms. These platforms themselves face limitations in producing unbiased, AI-ready datasets that accurately capture the intra- and inter-variation inherent in biological systems. Such limitations hinder the generalization capabilities of AI agents, which rely on comprehensive and high-quality data to function optimally. The absence of data from high-throughput techniques can lead to AI agents forming false hypotheses or causing harm. This risk is exacerbated when AI agents work with small, biased biological datasets, which may be affected by issues like batch effects.
Outlook
@@
Biomedical research is undergoing a transformative era with advances in computational intelligence. Presently, AI’s role is constrained to assistive tools in low-stake and narrow tasks where scientists can review the results. We outline agent-based AI to pave the way for systems capable of reflective learning and reasoning that consist of LLM-based systems and other ML tools, experimental platforms, humans, or even combinations of them. The continual nature of human-AI interaction and building trustworthy sandboxes [199], where AI agents can fail and learn from their mistakes, is one way to achieve this. This involves developing AI agents proficient in various tasks, such as planning discovery workflows with machine learning feedback loops for experiments and performing self-assessment to identify and seek out gaps in their knowledge, fostering natural and artificial intelligence.
Ensuring context-appropriate and user-specific agent behavior
To ensure agents behave as intended, it is essential to focus on their robustness and reliability by implementing evaluation protocols that test agents in diverse scenarios to identify potential vulnerabilities. Moreover, grounding agents in ethical guidelines and documentation, such as lab protocols and safety guidelines, is vital to align their actions with human values and safety standards. By addressing these aspects, we can ensure that the behavior of biomedical agents is both reliable and ethically compliant.
Concretely, we believe that in the early stages of technological adaptation, it is desirable to limit an agent’s capabilities to a subset of their full potential by restricting action spaces, thereby eliminating the chance of catastrophic risk (e.g., decisions resulting in loss of life). Similar precedents for technological adaptation are already in place for other autonomous systems with similar risk profiles, such as autonomous driving, where a staggered technological adaptation is motivated by ethical considerations.
Governance and responsible human-AI partnership
Managing errors requires designing strategies to diagnose, localize, and mitigate them. To diagnose errors internally, agents should use their reasoning abilities to build self-evaluation schemes, allowing them to assess their current status and actions. Externally, training independent anomaly detection and distribution shift models with domain knowledge of specific biomedical use cases can provide additional supervision to diagnose errors. Iterative agent interactions can result in cascading errors. To mitigate this, the evaluation agent can apply reverse reasoning chains to trace back to the initial error. Enhancing the adaptive reasoning abilities of agents is crucial for dynamically adjusting to changing conditions and rectifying errors as they occur.
In addressing the challenge of governance, it is our view that the required broad consensus is best reached in multi-disciplinary, cross-partisan, non-profit, and public institutions with the objective of public good. In this regard, we welcome the recent establishment of several public AI-focused safety institutions to facilitate that discussion. Our concrete advice for biomedical AI agents would be to establish focus groups or select committees with the required expertise that can define necessary ethical and technical evaluation standards, based on which concrete regulation can be determined (e.g., a necessary degree of human oversight and accountability structures). We furthermore advocate for standards and policies to be developed in the broadest international institutions possible to reduce the chance of risks simply being outsourced to jurisdictions with no existing or unenforceable regulations.
By fostering responsible human-AI partnerships and robust governance frameworks, we can harness the transformative potential of AI agents in biomedical research. This collaborative approach can pave the way for groundbreaking advances, eventually enhancing human health and well-being.
@@ Acknowledgments
We gratefully acknowledge the support of NIH R01-HD108794, NSF CAREER 2339524, US DoD FA8702-15-D-0001, awards from Harvard Data Science Initiative, Amazon Faculty Research, Google Research Scholar Program, AstraZeneca Research, Roche Alliance with Distinguished Scientists, Sanofi iDEA-iTECH Award, Pfizer Research, Chan Zuckerberg Initiative, John and Virginia Kaneb Fellowship award at Harvard Medical School, Aligning Science Across Parkinson’s (ASAP) Initiative, Biswas Computational Biology Initiative in partnership with the Milken Institute, and Kempner Institute for the Study of Natural and Artificial Intelligence at Harvard University. A.F. is supported by the Kempner Institute Graduate Fellowship. A.N. is supported by the Herchel Smith-Harvard Undergraduate Science Fellowship, the Yun Family Research Fellows Fund for Revolutionary Thinking, and the Summer Institute in Biomedical Informatics at Harvard Medical School. V.G. is supported by the Medical Research Council, MR/W00710X/1. Y.E. is supported by grant T32 HG002295 from the National Human Genome Research Institute and the NSDEG fellowship. The authors would like to thank Owen Queen, Alejandro Velez-Arce, and Ruth Johnson for their constructive comments on the draft manuscript. Any opinions, findings, conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the funders.
@@
Figure 1: Empowering biomedical research with AI agents. AI agents pave the way for "AI scientists" capable of skeptical learning and reasoning. These multi-agent systems consist of agents based on conversable large language models (LLMs) and can coordinate machine learning (ML) tools, experimental platforms, humans, or even combinations of them. Robotic agent, AI agent that operates robotic hardware for physical experiments; Database agent, AI agent that can information in databases via ‘function calling’ and APIs; Reasoning agent, AI agent capable of direct reasoning and reasoning with feedback; Hypothesis agent, AI agent that is creative and reflective when developing hypotheses, capable of characterizing its own uncertainty and using that as a driver to refine its scientific knowledge bases; Brainstorming agent, AI agent that generates a broad spectrum of research ideas; Search engine agent, AI agent that uses search engines as tools to rapidly gather information; Analysis agent, AI agent capable of analyzing experimental results to summarize findings and synthesize concepts; Experimental planning agent, AI agent that optimizes an experimental protocol for execution.
@@
Figure 2: Evolving use of data-driven models in research. Data-driven approaches, from databases and search engines, machine learning, and interactive learning models to advanced agent systems (Section Evolving use of data-driven models in biomedical research), have reshaped biomedical research throughout the last several decades. Dashed boxes represent studies focused predominantly on algorithmic machine learning innovation; solid-line boxes represent studies focused predominantly on biomedical discovery.
@@
Figure 3: Diverse configurations of AI agents in biomedicine – from an LLM-based AI agent to a multi-agent system with AI models, tools, and integrated physical devices. a. By programming an LLM with the role, one LLM-based agent, equipped with memory and reasoning abilities, performs multi-modal perception and utilizes a range of tools, e.g., web lab tools, to accomplish specified tasks. b-e. Leveraging AI agents equipped with diverse roles, perception modules, tools, and domain knowledge enables collaboration between agents and scientists. This collaboration can adopt various schemes, such as expert consultation, debate, brainstorming, and round table discussions. f. Multi-agent systems can establish a self-driving laboratory wherein numerous agents collaborate on multiple iterations of biological research assisted by humans. Each cycle of research encompasses the generation of hypotheses, the design of experiments, the execution of experiments both in silico and in vitro, and the analysis of results. Computing agent, AI agent that utilizes computational models as tools; Decision agent, AI agent that makes decisions in response to given conditions; Database agent, AI agent that retrieves relevant information from databases; Reasoning agent, AI agent capable of direct reasoning and reasoning with feedback; Expert agent, AI agent that provides professional consultation based on reliable sources, such as domain expertise, feedback from human experts, and the results of specific tools. Hypothesis agent, AI agent capable of reflective learning and reasoning to generate hypotheses; Planner agent, AI agent that devises plans for future actions; In silico/vitro agent, AI agent that uses tools in silico or in vitro environment.
@@
Figure 4: Four key modules of biomedical AI agents: perception, interaction, reasoning, and memory modules. Perception interprets multi-modal environmental data. Interaction facilitates engagement with the environment, encompassing human-agent interactions, multi-agent interactions, and tool use. Memory is responsible for the storage and retrieval of knowledge, while Learning focuses on the acquisition and updating of knowledge. Reasoning, with or without environmental feedback, plays a crucial role in planning and decision-making processes. Cross-modal alignment is a key technique for the perception of LLM-based agents, where inputs from different modalities are aligned within a text-centered representation space. This alignment enables the LLM to perceive and process various input modalities. Reasoning patterns for AI agents indicate transitions between reasoning thoughts. For instance, agents with a chain of thought pattern generate reasoning in a step-by-step manner.
@@
Figure 5: Illustration of components in biomedical AI agents. a. Use of a short-term memory module to recall previous relevant experiments for small molecule inhibitor design. b. Use of a long-term memory module to retrieve relevant information for target selection for a disease. c. Use of reasoning without scientist feedback in gene prioritization for phenotype analysis. d. Use of reasoning with feedback from scientists to select an alternative experimental approach. e. Combining perception, interaction, memory, and reasoning modules to study the selection against pathogenic mitochondria DNA in the germline.
@@
Figure 6: Challenges for AI agents in biomedical discovery. a. Shown are critical challenges–including robustness and reliability, evaluation protocols, dataset generation, governance, and risks—alongside b. strategic approaches to address them.
Autonomy Biomedical discovery Scientist-AI agent roles
levels Hypothesis Experiment Reasoning
@@
Autonomy levels Genetics (mutational effect modeling) Cell biology (drug resistance) Chemical biology (binder design)
[…]
Table 2: Examples of levels of autonomy of AI agents in genetics, cell biology, and chemical biology.
@@
In-context learning Ability to perform new tasks based on a handful of examples provided within the contextual prompt, without requiring explicit model training
Retrieval-augmented generation Techniques that make generative models to produce contextually relevant text by retrieving pertinent information and using it to inform the generation process